A photon and an electron each have an energy of (6.0 X 10^3 eV). What are their
wavelengths? Which of these would you use to probe atomic structures?
1
Expert's answer
2020-04-13T10:05:08-0400

Energy = 6000 eV


for photon

"\lambda=\frac{hc}{E}" "=\frac{1240}{E}nm=\frac{1240}{6000}nm=0.2067nm"


for electron

"\lambda=\frac{h}{\sqrt{2mE}}=\frac{6.63\times10^{-34}}{\sqrt{2\times9.1\times10^{-31}\times6000\times1.6\times10^{-19}}}=\frac{6.63\times10^{-34}}{41.8\times10^{-24}}=1.59\times10^{-9}m=1.59nm"


I would use electron to probe atomic structure